are the following lines perpendicular? y=6x-3 y=-1/6x-2
Answer by Cromlix(4381)

Lines that are perpendicular to one
another have slopes that multiply
together to give -1
y = 6x - 3
Slope = 6
y = -1/6x - 2
Slope = -1/6
6 * -1/6 = -1
Yes they are perpendicular to each other.
Hope this helps.
:-)